You throw a 20-N rock vertically into the air from ground level. You observe that when it is a height 15.9 m above the ground, it is traveling at a speed of 24.8 m/s upward.

A.Use the work-energy theorem to find its speed just as it left the ground.

B.Use the work-energy theorem to find its maximum height.
Answer #1

Gravitational acceleration = g = 9.81 m/s2

Weight of the rock = W = 20 N

Mass of the rock = m

Speed of the rock just as it left the ground = V0

Speed of the rock at a height 15.9 m above the ground = V1 = 24.8 m/s

Height of the rock where the speed of the rock is 24.8 m/s = H1 = 15.9 m

The potential energy and the kinetic energy of the rock on the ground is equal to the potential energy and kinetic energy of the rock at any height.

mV02/2 = mgH1 + mV12/2

V02 = 2gH1 + V12

V 2gH1 V2

Vo (2)(9.81) (15.9)(24.8)2

V0 = 30.446 m/s

Maximum height reached by the rock = H2

By conservation of energy the potential energy of the rock at maximum height is equal to the potential energy and kinetic energy of the rock when it is at 15.9 m above the ground.

mgH2 = mgH1 + mV12/2

На — Ні + 2g

(24.8)2 H2= 15.92)(9.81) (2) (9.81)

H2 = 47.247 m

A) Speed of the rock just as it left the ground = 30.446 m/s

B) Maximum height reached by the rock = 47.247 m